Larry wants to start an IRA that will have $420,000 in it when he retires in 29 years. How much should he invest semiannually in his IRA to do this if the interest is 6% compounded semiannually? A- $2354.65 B- $9288.16 C- $2767.16 D- $2754.16 E- $888.61
Answer #1
Let the investment be $x semi anually
=> x(1+0.03)^58+ x(1+0.03)^57...............x = 420000
=>x( 1.03^58+ 1.03^57...............1) = 420000
=>x(1.03^58-1)/(1.03-1) = 420000
=>x = 2767.16